A coin sold at auction in 2017 for $1,965,500. The coin had a face value of $5 when it was issued in 1794 and had previously been sold for $305,000 in 1971. a. At what annual rate did the coin appreciate from its first minting to the 1971 sale? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) What annual rate did the 1971 buyer earn on his purchase? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) At what annual rate did the coin appreciate from its first minting to the 2017 sale? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g. 32.16.) b. c. a. Rate of return from 1794 to 1971 b. Rate of return from 1971 to 2017 c. Rate of return from 1794 to 2017
Answer #1

a. Rate of return from 1794 to 1971 = Present value / Base value = (305000 / 5 ) * 100 = 6100000%

b. Rate of return from 1971 to 2017 = Present value / Base value = (1965500 / 305000) * 100 = 644.43%

c. Rate of return from 1794 to 2017 = Present value / Base value = (1965500 / 5 ) * 100 = 39310000%